# Distinguishing Distance Duality breaking models using electromagnetic and gravitational waves measurements

### v1

@@ 3.2 Supernovae
For our analysis, we simulate future SN data from the Legacy Survey of Space and Time [^1] https://rubinobservatory.org/for-scientists/rubin-101/key-numbers (LSST) [41]. We simulate $N_{\rm SN}=8800$ events, that corresponds to LSST expected number of events observed in the redshift range $z=0.1-1.0$, each with an error given by [58, 59]
@@ 3.3 Standard Sirens
The last observable that we include are GW, and in particular the Standard Sirens [60, 61]. GW represent a perturbation of spacetime that, in GR, propagates according to the equation
@@
where $h_{+,\times}(\tau,k)$ is the strain amplitude for the two polarization modes +,×+$\times$+ , ×, and $\tau$ is the conformal time. It can be demonstrated [62] that the strain amplitude $h_{+,\times}$ is inversely proportional to $d_{L}^{\rm GW}(z)$, meaning that the GW strain decreases with distance from the source as
@@
By accurately modelling the GW waveform, which depends on the masses, the orientation of the merging system, the spins of the objects and the distance, one can infer $d_{L}^{\rm GW}$ directly from the observed amplitude without relying on the traditional cosmic distance ladder. If an EM counterpart or an identified host galaxy also gives information about the redshift $z$, combining $d_{L}^{\rm GW}$ and $z$ provides a direct measurement of $H(z)$. These are known as Standard Sirens [61].
We simulate standard siren events by focusing on the merger of binary Neutron Stars (BNS). Our simulations are based on the expectations for the Einstein Telescope (ET) [39], a forthcoming third-generation ground-based GW detector [63]. We generate $\rm N=20000$ GW events, also known as injected signals, from BNS mergers with a rate modelled as [64, 36]
@@
For the position in the sky we assume that $\rm sin\theta_{\nu}$, where $\theta_{\nu}$ is the inclination angle, is uniformly distributed in the range $[-1,1]$; while for the redshift of the events we use a distribution
@@
The events are simulated exploiting a modified version of the darksirens [^2] https://gitlab.com/matmartinelli/darksirens code [65], which implements the merger rate of Eq. (3.7) and a monochromatic mass for the neutron stars of $1.4\ M_{\odot}$.
We compute the luminosity distance of a GW event using Eq. (2.6) and Eq. (2.7). The uncertainties on the luminosity distance $d_{L}^{\rm GW}$ and the detectability of the signals are evaluated using the public GWFish [^3] https://github.com/janosch314/GWFish code [66]. For what concerns the detectability we perform first a cut in the data based on the SNR ratio, considering an event to be detected if its signal-to-noise ratio (SNR) is larger than eight ($\rm SNR>8$), and then we analyze the EM counterpart detectability.
For the EM counterparts, we expect mainly two types of emission: a Gamma Ray Burst (GRB) and a Kilonova [67]. Following the merger, a highly collimated and relativistic jet is formed, which emits the GRB prompt emission and its afterglow. The prompt emission is detected right after the merger, at high (gamma-ray) energy, and with a duration below 2 sec. Once the jet reaches the interstellar medium, it starts to decelerate and spread, producing the afterglow. The latter is mainly synchrotron emission spanning the whole spectrum from X-rays to radio wavelengths and can last even years after the merger.
A small amount of fast-moving neutron-rich ejecta emits an isotropic thermal emission, the kilonova, peaking in the infrared. A larger mass neutron-free wind along the polar axis produces the kilonova emission peaking at optical wavelenghts.
@@
In this work, we focus on the afterglow emission and we assume that each GW event produces a GRB (even if this is not necessarily true [68]). In particular, we use the luminosity distance and the inclination from the GW counterparts to generate the afterglow flux using the code afterglowpy [^4] https://github.com/geoffryan/afterglowpy [69]. The flux depends also on the energetics and the mycrophysics of the event, which we fix to the one estimated in population studies of short GRBs [70, 71]. In particular, we assume a Gaussian structured jet and we fix the jet opening angle to 6o, the isotropic-equivalent kinetic energy $E_{0}$ to $4\times 10^{51}$ erg, the circum-burst density $n_{0}$ to $6.4\times 10^{-4}$ cm-3, the fraction of energy in the magnetic field $\epsilon_{B}$ and in the electrons $\epsilon_{e}$ both to 0.1, and the slope of the electron distribution $p$ to 2.3.
As a reference we use the sensitivity of the Vera Rubin Observatory (LSST), that will become operational at the end of 2025. It will reach a sensitivity of $24.5$ magnitudes in the R band [41], so a flux threshold of about $F_{\rm th}=5.8\times 10^{-4}\,\,\rm mJy$. Taking into consideration all these characteristics we obtain that around $40\%$ of the GW events detected would have an EM counterpart, this can be seen in Figure 3. As expected, the smaller is the inclination, the higher is the chance to detect the source.
Figure 3: Detected GW events as a function of redshift and inclination angle $\theta_{\nu}$, along with their corresponding peak flux. The events are simulated assuming a fiducial cosmology from Pantheon+SH0ES [72], with $H_{0}=73.4$ and $\Omega_{m}=0.306$. The grey points represent events detected by ET that survive the SNR cut, while the coloured points indicate events that would also be detected with an associated electromagnetic counterpart.
@@ 4.1 Reconstructing the posterior distribution
In order to constrain possible deviations from the standard DDR, we use statistical methods to obtain constraints on the $\eta(z)$ function, which we parameterize following the discussion of section 2. This process is done by reconstructing the posterior distribution of $\epsilon$, together with other cosmological parameters that will be discussed in subsection 4.2. This reconstruction is grounded in Bayesian statistics [73], with the comparison between data and theoretical predictions handled by the public likelihood code Cobaya [^5] https://github.com/CobayaSampler/cobaya [74], for which we construct external likelihood modules able to handle the datasets discussed in section 3.
For what concerns the theoretical modelling, we obtain the standard cosmological quantities through CAMB [^6] https://camb.readthedocs.io/ [75, 76], interfaced with a theory code that includes the expression for $\eta$ and the modifications to $d_{L}$. This setup  [^7] The code will be publicly available upon publication of the paper in the journal. allows Cobaya to perform Markov Chain Monte Carlo (MCMC) sampling to reconstruct the posterior distribution [77, 78].
@@ 4.2 Models and sampled parameters
  • Both SN and GW are consistent with the DDR and no violation happens ($\Lambda$CDM Universe);
  • Both SN and GW violate the DDR, and this could be addressed to some modifications in the theory of gravity (full-breaking Universe);
  • Only SN violate the DDR and this would be the case of dust or decay of photons into non-standard model particles (EM-breaking Universe).
We generate simulated observations corresponding to these scenarios, using the fiducial values from Pantheon+SH0ES [72] shown in Table 1. We analyze these simulated observations sampling the DDR violations parameters, together with the cosmological ones, assuming the prior distributions reported in Table 2.
@@
All the sampled parameters are assumed to have a uniform prior distribution informed by [79] results or [80]. We stress the fact that no Gaussian priors have been applied to $\Omega_{b}h^{2}$ for BAO data or to $M_{B}$ for SN data. This implies that when considering only BAO and SN data, without including GW, the sampling is performed on $r_{d}$ for BAO and by marginalizing over $M_{B}$ for SN. However, when GW are included, they provide independent information on $H_{0}$, allowing us to sample on $\Omega_{b}h^{2}$ for BAO while maintaining a flat prior, and to sample over both $M_{B}$ and $H_{0}$ for SN, always using uniform priors.
@@ 5.1 Universal Violation Scenario - UVS
Figure 4: (a) : Error associated with the UVS parameter $\epsilon_{0}$ for different combinations of observables, shown for both the $\Lambda$CDM case (solid blue) and the breaking case (dashed pink). (b) : Constraints in the $\epsilon_{0}$$\Omega_{m}$ parameter space obtained assuming a universal violation scenario (UVS) affecting both the electromagnetic and gravitational sectors. If the true physical behavior instead involves a violation only in the electromagnetic sector, applying the UVS framework can artificially induce a tension in $\epsilon_{0}$ between the BAO+SN and BAO+GW datasets.
Another interesting result that highlights the impact of the inclusion of both EM and gravitational observables, is shown in 4(b), where we present the constraints in the $(\epsilon_{0},\Omega_{m})$ parameter space obtained under the UVS assumption. We can notice that, if this is not the correct physical framework to describe the actual behaviour of the system (i.e. in the EM-breaking Universe), applying such an incorrect model can induce a tension in the UVS parameter $\epsilon_{0}$ between the BAO+SN and BAO+GW datasets. This outcome demonstrates how the use of an incorrect theoretical model can lead to misleading discrepancies between observational probes, emphasising the importance of employing a physically accurate framework when interpreting cosmological data. Furthermore, this highlights the complementarity of SN and GW, as the individual use of any of these two probes would not allow us to notice the inaccurate modelling of the physical effect.
The results concerning this case, both in the $\Lambda$CDM, EM-breaking and full-breaking Universes, are shown in Table 3.
Table 3: Constraints on $\Omega_{m}$ and $\epsilon_{0}$ for different dataset combinations, comparing the $\Lambda$CDM Universe and the full-breaking Universe within the UVS assumption. For the EM-breaking case the joint analysis of SN+GW is not computed, since they are in tension.
@@
$\epsilon_{0}$ $-0.0070\pm 0.0074$ $0.3011\pm 0.0040$
UVS - Full-breaking $\Omega_{m}$ $0.329^{+0.084}_{-0.069}$ $0.3014\pm 0.0040$ $0.1038\pm 0.0059$ $0.3015\pm 0.0040$
$\epsilon_{0}$ $0.113^{+0.077}_{-0.047}$ $0.0946\pm 0.0065$ $0.0950\pm 0.0051$ $0.0950\pm 0.0047$
@@ 5.2 Independent Violation Scenario
In the IVS, we instead allow for deviations from the standard DDR to occur independently in the electromagnetic and gravitational sectors. Accordingly, the analysis introduces two separate parameters: $\epsilon_{\rm EM}$ and $\epsilon_{\rm GW}$, which capture possible violations in each sector. The sampled parameters and priors for this analysis are listed in Table 2. We explore the three distinct cases for which we simulated the observations: the $\Lambda$CDM, full-breaking and EM-breaking Universes.
These three cases are illustrated in Figure 5, and they yield clearly distinguishable patterns in the parameter space. This result highlights the power of combining SN and GW data with BAO measurements to effectively constrain deviations from the standard $\Lambda$CDM cosmology and the validity of the Etherington relation. Moreover, this approach allows us to distinguish the physical origin of such deviations. The addition of GW measurements is essential in this context. When the analysis is limited to the electromagnetic sector alone, the scenarios involving purely electromagnetic violations (pink) and combined electromagnetic and gravitational violations (blue) become indistinguishable, despite reflecting fundamentally different physics. Without access to the $\epsilon_{0}^{\rm GW}$ axis, the blue case would effectively collapse onto the pink one in the $\epsilon_{\rm EM}$$\Omega_{m}$ parameter space, preventing us from distinguishing between the two scenarios. However, making this distinction is crucial, as they correspond to fundamentally different physical contexts. In particular, the pink case, which involves a violation only in the electromagnetic sector, could point to mechanisms such as photon absorption by intergalactic dust or exotic particle physics processes. In contrast, the blue case suggests a violation that also affects the gravitational sector, potentially arising in modified gravity frameworks. A summary of the constraint obtained with this configuration in the three different physical scenarios can be seen in Table 4.
Figure 5: Constraints in the $\epsilon_{0}^{\rm EM}$$\epsilon_{0}^{\rm GW}$ parameter space for three different physical scenarios: $\Lambda$CDM (orange), violation in the electromagnetic sector only (pink), and violation in both the electromagnetic and gravitational sectors (blue). The black dashed line represents the case $\epsilon_{0}^{\rm EM}=\epsilon_{0}^{\rm GW}$.
@@
$\Lambda$CDM $0.3012\pm 0.0041$ $-0.0004\pm 0.0050$ $-0.0068\pm 0.0070$
$\epsilon_{\mathrm{EM}}$ breaking $0.3011\pm 0.0040$ $0.0950\pm 0.0052$ $-0.0068\pm 0.0069$
$\epsilon$ breaking $0.3016\pm 0.0040$ $0.0954\pm 0.0052$ $0.0948\pm 0.0066$
@@ 5.3 Padé parametrization
The final results we present concern the Padé parametrization introduced in Eq. (2.16). This is a phenomenological approximation to hypothetical models in which the mechanism responsible for deviations from the standard DDR is naturally suppressed at high redshift. Such an approximation allows to avoid the modelling of early time physics in a context where DDR is violated, which could lead to complicated theoretical modelling. In this analysis, we explore the consequences of using such an approximation, when instead the DDR breaking mechanism is instead active at all redshifts.
@@
This comparison is performed within both the UVS and IVS frameworks by generating mock data using the constant $\epsilon$ parametrization defined in Eq. (2.7) and then analyzing this data assuming either the Padé or constant $\epsilon$ model. This is shown in Figure 6. In each figure, we compare the same dataset analyzed under both assumptions for $\epsilon$, presenting only the two-dimensional contours from the joint BAO+SN+GW analysis. Across all breaking scenarios, UVS, IVS-EM, and IVS-full, we observe a shift in the inferred DDR parameters when the wrong model is used. However, this shift is not statistically significant, as the results remain consistent within the 95%. Consequently, we can conclude that the Padé approximation can be safely used, at least for the redshift range of interest of the observables used in this work.
@@
UVS $\epsilon$ const $0.3015\pm 0.0040$ $0.0950\pm 0.0047$
UVS Padé $0.3011\pm 0.0041$ $0.0892\pm 0.0048$
IVS-EM $\epsilon$ const $0.3011\pm 0.0040$ $0.0950\pm 0.0052$ $-0.0068\pm 0.0069$
IVS-EM Padé $0.3014\pm 0.0041$ $0.0927\pm 0.0052$ $-0.0063\pm 0.0070$
IVS-EM+GW $\epsilon$ const $0.3016\pm 0.0040$ $0.0954\pm 0.0052$ $0.0948\pm 0.0066$
IVS-EM+GW Padé $0.3013\pm 0.0041$ $0.0924\pm 0.0052$ $0.0789\pm 0.0078$

### v2

@@ 3.2 Supernovae
For our analysis, we simulate future SN data from the Legacy Survey of Space and Time [^1] https://rubinobservatory.org/for-scientists/rubin-101/key-numbers (LSST) [Bianco2022]. We simulate $N_{\rm SN}=8800$ events, that corresponds to LSST expected number of events observed in the redshift range z=0.11.0z=0.1-1.0, each with an error given by [2010ApJ...709.1420G, Astier:2014swa]
@@ 3.3 Standard Sirens
The last observable that we include are GW, and in particular the Standard Sirens [Holz_2005, Nissanke2010]. GW represent a perturbation of spacetime that, in GR, propagates according to the equation
@@
where $h_{+,\times}(\tau,k)$ is the strain amplitude for the two polarization modes $\times$, and $\tau$ is the conformal time. It can be demonstrated [1986Natur.323..310S] that the strain amplitude $h_{+,\times}$ is inversely proportional to $d_{L}^{\rm GW}(z)$, meaning that the GW strain decreases with distance from the source as
@@
By accurately modelling the GW waveform, which depends on the masses, the orientation of the merging system, the spins of the objects and the distance, one can infer $d_{L}^{\rm GW}$ directly from the observed amplitude without relying on the traditional cosmic distance ladder. If an EM counterpart or an identified host galaxy also gives information about the redshift $z$, combining $d_{L}^{\rm GW}$ and $z$ provides a direct measurement of $H(z)$. These are known as Standard Sirens [Nissanke2010].
We simulate standard siren events by focusing on the merger of binary Neutron Stars (BNS). Our simulations are based on the expectations for the Einstein Telescope (ET) [abac2025scienceeinsteintelescope], a forthcoming third-generation ground-based GW detector [Branchesi_2023]. We assume a triangular configuration with 10 km arms, located in Sardinia. We generate $\rm N=20000$ GW events, also known as injected signals, from BNS mergers with a rate modelled as [Cutler:2009qv, Hogg:2020ktc]
@@
For the position in the sky we assume that $\rm sin\theta_{\nu}$, where $\theta_{\nu}$ is the inclination angle, is uniformly distributed in the range [1,1][-1,1]; while for the redshift of the events we use a distribution
@@
The events are simulated exploiting a modified version of the darksirens [^2] https://gitlab.com/matmartinelli/darksirens code [darksirens], which implements the merger rate of Eq. (3.7) and a monochromatic mass for the neutron stars of $1.4\ M_{\odot}$.
We compute the luminosity distance of a GW event using Eq. (2.6) and Eq. (2.7). The uncertainties on the luminosity distance $d_{L}^{\rm GW}$ and the detectability of the signals are evaluated using the public GWFish [^3] https://github.com/janosch314/GWFish code [Dupletsa_2023]. For what concerns the detectability, we first perform a cut in the data based on the SNR ratio, considering an event to be detected if its signal-to-noise ratio (SNR) is larger than eight ($\rm SNR>8$), and then we analyze the EM counterpart detectability. The number of detected BNS mergers increases if ET is in a network of third-generation GW detectors. As an example, ET plus one Cosmic Explorer [Reitze:2019iox] would increase of a factor $\sim 3$ the number of detected BNS, and a factor $\sim 4$ with two Cosmic Explorers. We also note that even changing the ET configuration (two L-shaped detectors instead of a triangle) can lead to an almost 3 times higher detection rate [Branchesi_2023]. Given all these uncertainties, we choose to focus only ET, with a triangular configuration.
For the EM counterparts, we expect mainly two types of emission: a Gamma Ray Burst (GRB) and a Kilonova [metzger2017welcomemultimessengereralessons]. Following the merger, a highly collimated and relativistic jet is formed, which emits the GRB prompt emission and its afterglow. The prompt emission is detected right after the merger, at high (gamma-ray) energy, and with a duration below 2 sec. Once the jet reaches the interstellar medium, it starts to decelerate and spread, producing the afterglow. The latter is mainly synchrotron emission spanning the whole spectrum from X-rays to radio wavelengths and can last for even years after the merger.
A small amount of fast-moving neutron-rich ejecta emits an isotropic thermal emission, the kilonova, peaking in the infrared. A larger mass of neutron-free wind along the polar axis produces the kilonova emission peaking at optical wavelengths.
@@
In this work, we focus on the afterglow emission and we assume that each GW event produces a GRB (even if this is not necessarily true [Metzger_counterpart]). In particular, we use the luminosity distance and the inclination from the GW counterparts to generate the afterglow flux using the code afterglowpy [^4] https://github.com/geoffryan/afterglowpy [Ryan_2020]. The flux also depends on the energetics and the mycrophysics of the event, which we fix to the one estimated in population studies of short GRBs [RoucoEscorial_2023, Fong_2015]. In particular, we assume a Gaussian structured jet and we fix the jet opening angle to 6o, the isotropic-equivalent kinetic energy $E_{0}$ to $4\times 10^{51}$ erg, the circum-burst density $n_{0}$ to $6.4\times 10^{-4}$ cm-3, the fraction of energy in the magnetic field $\epsilon_{B}$ and in the electrons $\epsilon_{e}$ both to 0.1, and the slope of the electron distribution $p$ to 2.3.
As a reference, we use the sensitivity of the Vera Rubin Observatory (LSST), which will become operational at the end of 2025. It will reach a sensitivity of $24.5$ magnitudes in the R band [Bianco2022], so a flux threshold of about $F_{\rm th}=5.8\times 10^{-4}\,\,\rm mJy$. Taking into consideration all these characteristics we obtain that around $40\%$ of the GW events detected would have an EM counterpart, this can be seen in Figure 3. As expected, the smaller the inclination, the higher the chance to detect the source.
Figure 3: Detected GW events as a function of redshift and inclination angle $\theta_{\nu}$, along with their corresponding peak flux. The events are simulated assuming a fiducial cosmology from Pantheon+SH0ES [Riess_2022], with $H_{0}=73.4$ and $\Omega_{m}=0.306$. The grey points represent events detected by ET that survive the SNR cut, while the coloured points indicate events that would also be detected with an associated electromagnetic counterpart.
@@ 4.1 Reconstructing the posterior distribution
In order to constrain possible deviations from the standard DDR, we use statistical methods to obtain constraints on the $\eta(z)$ function, which we parameterize following the discussion of section 2. This process is done by reconstructing the posterior distribution of $\epsilon$, together with other cosmological parameters that will be discussed in subsection 4.2. This reconstruction is grounded in Bayesian statistics [bayesian], with the comparison between data and theoretical predictions handled by the public likelihood code Cobaya [^5] https://github.com/CobayaSampler/cobaya [Cobaya], for which we construct external likelihood modules able to handle the datasets discussed in section 3.
For what concerns the theoretical modelling, we obtain the standard cosmological quantities through CAMB [^6] https://camb.readthedocs.io/ [Lewis:1999bs, 2012JCAP...04..027H], interfaced with a theory code that includes the expression for $\eta$ and the modifications to $d_{L}$. The setup used in this work, made publicly available as CANDI [^7] https://github.com/chiaradeleo1/CANDI [Lewis:1999bs, 2012JCAP...04..027H], allows Cobaya to perform Markov Chain Monte Carlo (MCMC) sampling to reconstruct the posterior distribution [Lewis:2002ah, Lewis:2013hha].
@@ 4.2 Models and sampled parameters
[…]
  • Both SN and GW violate the DDR, and this could be addressed by some modifications in the theory of gravity (full-breaking Universe);
  • Only SN violate the DDR, and this would be the case of dust or decay of photons into non-standard model particles (EM-breaking Universe).
We generate simulated observations corresponding to these scenarios, using the fiducial values from Pantheon+SH0ES [Riess_2022] shown in Table 1. We analyze these simulated observations sampling the DDR violations parameters, together with the cosmological ones, assuming the prior distributions reported in Table 2.
@@
All the sampled parameters are assumed to have a uniform prior distribution informed by [Planck18] results or [Brout_2022]. We stress the fact that no Gaussian priors have been applied to $\Omega_{b}h^{2}$ for BAO data or to $M_{B}$ for SN data. This implies that when considering only BAO and SN data, without including GW, the sampling is performed on $r_{d}$ for BAO and by marginalizing over $M_{B}$ for SN. However, when GW are included, they provide independent information on $H_{0}$, allowing us to sample on $\Omega_{b}h^{2}$ for BAO while maintaining a flat prior, and to sample over both $M_{B}$ and $H_{0}$ for SN, always using uniform priors.
@@ 5.1 Universal Violation Scenario - UVS
[…]
Another interesting result that highlights the impact of the inclusion of both EM and gravitational observables is shown in 4(b), where we present the constraints in the $(\epsilon_{0},\Omega_{m})$ parameter space obtained under the UVS assumption. We can notice that, if this is not the correct physical framework to describe the actual behaviour of the system (i.e., in the EM-breaking Universe), applying such an incorrect model can induce a tension in the UVS parameter $\epsilon_{0}$ between the BAO+SN and BAO+GW datasets. This outcome demonstrates how the use of an incorrect theoretical model can lead to misleading discrepancies between observational probes, emphasising the importance of employing a physically accurate framework when interpreting cosmological data. Furthermore, this highlights the complementarity of SN and GW, as the individual use of any of these two probes would not allow us to notice the inaccurate modelling of the physical effect.
[…]
@@
$\epsilon_{0}$ $0.0070\pm 0.0074$ $0.1038\pm 0.0059$
UVS - Full-breaking $\Omega_{m}$ $0.329^{+0.084}_{-0.069}$ $0.3014\pm 0.0040$ $0.3011\pm 0.0040$ $0.3015\pm 0.0040$
$\epsilon_{0}$ $113^{+0.077}_{-0.047} 0.0$$946\pm 0.0065 0.0$$950\pm 0.0051 0.0$$0950\pm 0.0047$
@@ 5.2 Independent Violation Scenario
[…]
@@
[…]
$\epsilon_{\mathrm{EM}}$ breaking $3011\pm 0.0040 0.0$$950\pm 0.0052 0.0$$0068\pm 0.0069$
$\epsilon$ breaking $3016\pm 0.0040 0.0$$954\pm 0.0052 0.0$$0948\pm 0.0066$
@@ 5.3 Padé parametrization
The final results we present concern the Padé parametrization introduced in Eq. (2.16). This is a phenomenological approximation to hypothetical models in which the mechanism responsible for deviations from the standard DDR is naturally suppressed at high redshift. Such an approximation allows to avoid the modelling of early time physics in a context where DDR is violated, which could lead to complicated theoretical modelling. In this analysis, we explore the consequences of using such an approximation, when the DDR breaking mechanism is instead active at all redshifts.
@@
This comparison is performed within both the UVS and IVS frameworks by generating mock data using the constant $\epsilon$ parametrization defined in Eq. (2.7) and then analyzing this data assuming either the Padé or constant $\epsilon$ model. This is shown in Figure 6. In each figure, we compare the same dataset analyzed under both assumptions for $\epsilon$, presenting only the two-dimensional contours from the joint BAO+SN+GW analysis. Across all breaking scenarios, UVS, IVS-EM, and IVS-full, we observe a shift in the inferred DDR parameters when the wrong model is used. However, this shift is not statistically significant, as the results remain consistent within the 95% confidence level. Consequently, we can conclude that the Padé approximation can be safely used, at least for the redshift range of interest of the observables used in this work.
@@
UVS $\epsilon$ const $3015\pm 0.0040 0.0$$0950\pm 0.0047$
UVS Padé $3011\pm 0.0041 0.0$$0892\pm 0.0048$
IVS-EM $\epsilon$ const $3011\pm 0.0040 0.0$$950\pm 0.0052 0.0$$0068\pm 0.0069$
IVS-EM Padé $3014\pm 0.0041 0.0$$927\pm 0.0052 0.0$$0063\pm 0.0070$
IVS-EM+GW $\epsilon$ const $3016\pm 0.0040 0.0$$954\pm 0.0052 0.0$$0948\pm 0.0066$
IVS-EM+GW Padé $3013\pm 0.0041 0.0$$924\pm 0.0052 0.0$$0789\pm 0.0078$